OpenStudy (anonymous):

complete pattern 50,43,36,29,__,__

OpenStudy (anonymous):

22

OpenStudy (anonymous):

thanks and ?

OpenStudy (anonymous):

22,15

OpenStudy (anonymous):

22,15

jimthompson5910 (jim_thompson5910):

You're subtracting 7 each time So the next two numbers are 29 - 7 = 22 and 22 - 7 = 15

OpenStudy (anonymous):

50-43=7 36-7=29 29-7=22 22-7=15